OpenStudy (kmullis6):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (kmullis6):

I don't get what it means by ''all real numbers'' /:

OpenStudy (kmullis6):

@tHe_FiZiCx99 @Mashy @beccaboo333

OpenStudy (anonymous):

-3(x+3) = -3(x+1) - 5 -> -3(x+3-x-1) = -5 -> -3(2) = -5, which is never true, so no solutions.

OpenStudy (the_fizicx99):

"All real numbers means, positive or negatives, large or small, whole numbers or decimal numbers are all Real Numbers." Apply distributive property to -3(x+3) and -3(x + 1)

OpenStudy (kmullis6):

If I'm doing it right.... -3(6+3)=-27 -3(6+1)=-21

OpenStudy (kmullis6):

I'm trying to find the same outcome right? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

No, don't use the options... solve it without them. -3(x + 3) = -3x - 9 -3(x + 1) = ?

OpenStudy (kmullis6):

So basically -3+1?

OpenStudy (the_fizicx99):